x^2+3x-40=0
How would you solve this equation?

Found 2 solutions by kapilsinghi, elima:
Answer by kapilsinghi(68)

x^2+3x-40=0
x^2+8x-5x-40=0
x(x+8)-5(x+8)=0
(x-5)(x+8)=0
x=5 or x=-8
kapilsinghi123@gmail.com

Answer by elima(1433)

x^2+3x-40=0
First step would be:
(x+8)(x-5)=0
THen solve each side for x;
x+8=0
x=-8
x-5=0
x=5
So we have;
x=-8, or x=5
=)